Sulfur trioxide is made from the reaction of sulfur dioxide and oxygen. How many grams of SO3 can be made from the reaction of 1.91 g SO2 with 1.29 g O2?
Answer #1

Ans :

The balanced reaction is given as :

2SO2 + O2 = 2SO3

Number of mol of SO2 = mass / molar mass

= 1.91 g / 64.0638 g/mol = 0.0298 mol

Number of mol of O2 = 1.29 g / 31.999 g/mol = 0.0403 mol

Each mol O2 needs 2 mol SO2

So here , number of mol of SO2 required = 2 x 0.0403 mol = 0.0806 mol

Since we have only 0.0298 mol of SO2 , so SO2 is limiting reactant here

Then the number of mol of SO3 formed = mol x molar mass

= 0.0298 mol x 80.0632 g/mol

= 2.39 g